# ProtoSiTex: Learning Semi-Interpretable Prototypes for Multi-label Text Classification

### v1

@@ 3.1. ProtoSiTex Architecture
For sentence-level ground truth $\mathcal{G}_{2}$, we aggregate sub-sentence labels via $\mathcal{M}^{2}$. Document-level ground truth $\mathcal{G}_{3}\in\mathbb{R}^{1\times l}$ is directly taken from the dataset. It is worth noting that prototype and sentence-level labels are generated dynamically and depend on the current prototype configuration during the clustering phase.
3.2. Training Strategy of ProtoSiTex
We adopt an alternating dual-phase training strategy, where each epoch comprises $i_{1}$ clustering and $i_{2}$ classification iterations. Prototype parameters are updated in the clustering phase (with classifiers frozen), and vice versa in the classification phase. This decoupling stabilizes training, prevents early overfitting, and allows prototypes to evolve independently before being used for label prediction. We now describe the loss functions associated with each phase.
3.2.1. Loss functions for Clustering Phase
The clustering loss comprises two complementary components: proximity-based losses, which enforce alignment between sub-sentence embeddings and prototypes, and structure regularization losses, which ensure sparsity and diversity among the learned prototypes.
(i) Proximity-based Losses: These losses promote semantic alignment and coverage. The prototype matching loss ensures that each sub-sentence $\mathcal{E}_{i}$ is closely aligned to its nearest prototype $\mathcal{P}_{j^{*}_{i}}$, encouraging relevance. Simultaneously, the interpretability loss ensures each prototype remains active by encouraging proximity to at least one sub-sentence, thus achieving coverage and preventing prototype collapse. The joint formulation is given by:
(5) $\mathcal{L}_{prox}=-\frac{1}{t}\sum_{i=1}^{t}\left(\frac{\mathcal{E}_{i}^{\top}\mathcal{P}_{j^{*}_{i}}}{\lVert\mathcal{E}_{i}\rVert\lVert\mathcal{P}_{j^{*}_{i}}\rVert}\right)+\frac{1}{q}\sum_{j=1}^{q}\left(1-\frac{\mathcal{E}_{i}^{\top}\mathcal{P}_{j^{*}_{i}}}{\lVert\mathcal{E}_{i}\rVert\lVert\mathcal{P}_{j^{*}_{i}}\rVert}\right)$
were, $j^{*}_{i}$ is the index of the prototype closest to sub-sentence $\mathcal{E}_{i}$ based on cosine similarity.
(ii) Structure Regularization Losses: To enhance interpretability, we impose sparsity and diversity through regularization. The sparsity loss encourages selective activation, ensuring a limited set of sub-sentences influences each prototype. This reduces overlap and enhances clarity. The diversity loss, implemented via orthogonality regularization, pushes prototypes apart in embedding space, minimizing redundancy. The combined regularization is:
(6) $\mathcal{L}_{reg}=-\frac{1}{q}\sum_{j=1}^{q}\log\left(\sum_{i=1}^{t}\frac{\exp(-\left\lVert\mathcal{E}_{i}-\mathcal{P}_{j}\right\rVert_{2})}{\sum_{k=1}^{q}\exp(-\left\lVert\mathcal{E}_{i}-\mathcal{P}_{k}\right\rVert_{2})}+\epsilon\right)+\frac{1}{q(q-1)}\sum_{\begin{subarray}{c}i,j=1;\\ i\neq j\end{subarray}}^{q}\left(\frac{\mathcal{P}_{i}^{\top}\mathcal{P}_{j}}{\lVert\mathcal{P}_{i}\rVert\lVert\mathcal{P}_{j}\rVert}\right)^{2}$
The first term encourages peaked assignment distributions (sparsity), and the second term enforces pairwise dissimilarity (diversity) between prototypes.
To enforce semantic alignment with classes, we augment the clustering loss with the supervised classification loss $\mathcal{L}_{cs}$ (Eq. 8). The final clustering-phase loss ($\mathcal{L}_{cl}$) is:
(7) $\scriptsize\mathcal{L}_{cl}=\alpha_{1}~\mathcal{L}_{prox}+\alpha_{2}~\mathcal{L}_{reg}+\alpha_{3}~\mathcal{L}_{cs}$
Here, $\alpha_{1},\alpha_{2},\alpha_{3}$ are tunable weights summing to 1, controlling the influence of each component.
3.2.2. Loss functions for Classification Phase
The classification phase is designed to facilitate hierarchical supervision by capturing semantic patterns from fine-grained (sub-sentence) to coarse-grained (document) levels. To train this phase, we use three loss components: (i) a prototype-to-label interpretability loss that reinforces the semantic alignment between prototypes and classes, (ii) a sentence-wise supervision loss that captures mid-level contextual cues, and (iii) a document-level loss leveraging the provided annotations. Each loss component is computed using binary cross entropy (Paszke et al., 2019), and the combined loss is expressed as:
(8) $\scriptsize\mathcal{L}_{cs}=\lambda_{1}~\mathcal{L}_{1}(\mathcal{G}_{1},\tilde{\mathcal{G}}_{1})+\lambda_{2}~\mathcal{L}_{2}(\mathcal{G}_{2},\tilde{\mathcal{G}}_{2})+\lambda_{3}~\mathcal{L}_{3}(\mathcal{G}_{3},\tilde{\mathcal{G}}_{3})$
Here, $\mathcal{G}_{i}$ and $\tilde{\mathcal{G}}_{i}$ (for $i=1,2,3$) represent ground-truth and predicted labels at prototype, sentence, and document levels. The weights $\lambda_{i}$ control each loss term, with $\sum_{i=1}^{3}\lambda_{i}=1$. Sub-sentence and prototype-level annotations are optional; when unavailable, $\lambda_{1}$ and $\lambda_{2}$ can be reduced or set to zero, enabling the model to rely solely on document-level labels and support flexible supervision.
This phase refines prototypes into interpretable classes, enhancing clustering and enabling generalizable, multi-granular predictions. Next, we describe the prototype-based inference process.
3.3. Prototype Projection and Label Classification
@@
Each prototype $p_{i}\in\mathcal{P}$ is a high-dimensional vector residing in the same semantic space as sub-sentence embeddings. To ensure interpretability, we associate each prototype with its top-$K$ most similar sub-sentences from the training set, using cosine similarity. The top-$K$ nearest sub-sentences $\mathcal{T}_{K}(p_{i})$ are consolidated and passed to Gemini (Team et al., 2023) via an API to generate a concise summary reflecting their shared theme. This strategy ensures each prototype is grounded in semantically diverse and representative text snippets, offering robust explanations.
@@
Given a test document $\mathcal{D}$, we segment it into sub-sentences $\{s_{ij}\}$, and encode them using the LM-based encoder to get $\mathcal{E}\in\mathbb{R}^{t\times d}$. These embeddings are projected to the prototype space via the trained multi-head attention module to produce prototype-aware representations $\mathcal{H}$, which are then passed through the proto-to-label classifier $\tilde{\mathcal{G}}_{1}$. Sentence- and document-level predictions ($\tilde{\mathcal{G}}_{2}$, $\tilde{\mathcal{G}}_{3}$) are obtained via hierarchical mappings using $\mathcal{M}^{2}$ and $\mathcal{M}^{1}$. The final class labels for the document are taken as $\tilde{\mathcal{G}}_{3}$. Simultaneously, we generate textual explanations by linking each sub-sentence to its closest prototype, and retrieving that prototype’s associated exemplars $\mathcal{T}_{K}(p_{i})$. This enables transparent reasoning behind predictions, making ProtoSiTex both accurate and interpretable.
@@
ProtoSiTex is best described as semi-interpretable because it combines prototype-based explanations with deep neural components. On the one hand, its adaptive prototypes provide human-understandable evidence by aligning sub-sentences with semantically coherent exemplars, allowing users to see why particular text spans support specific class predictions. On the other hand, the model relies on latent mechanisms, multi-head attention, embedding transformations, and hierarchical aggregation across sub-sentence, sentence, and document levels, which are not directly transparent to human reasoning. As a result, the explanations are faithful and intuitive at the prototype level but do not offer full interpretability of the entire decision process. This balance between interpretability and accuracy justifies labeling the model as semi-interpretable rather than fully interpretable.
4. Experiments
All experiments were conducted on a Dell Precision 5860 Tower equipped with PyTorch 2.5.1, CUDA 12.2, an NVIDIA RTX A6000 GPU with 48 GB VRAM, and 256 GB RAM. We employed RoBERTa-Large (Liu et al., 2019) as the backbone for encoding subsentences in our framework. The HR dataset was split into training, validation, and test sets in a 7:2:17{:}2{:}1 ratio. For IMDb and TweetEVAL, we used the given training/ validation/ test sets. Each experiment was run five times, and the mean test performance is reported. Additional model configuration details are provided in the Appendix A.
4.1. Datasets
@@
To evaluate the effectiveness of ProtoSiTex and compare it against state-of-the-art (SOTA) methods, we conducted comprehensive experiments on benchmark datasets IMDb (Maas et al., 2011), TweetEVAL (Barbieri et al., 2020), and our newly proposed Hotel Reviews (HR) dataset. These datasets encompass a diverse range of classification tasks, binary, multi-class, and multi-label, respectively.
@@
(i) IMDb (Maas et al., 2011): Consisting of 50K movie reviews for binary sentiment classification (positive/ negative).
@@
(ii) TweetEVAL (Barbieri et al., 2020): A multiclass dataset comprising 5052 tweets, with each tweet annotated with one of four emotion labels: anger, joy, optimism, or sadness.
@@
(iii) Proposed Dataset: Hotel Reviews (HR): To support fine-grained, interpretable, multi-aspect classification in the hospitality domain, we introduce a novel dataset comprising hotel reviews, curated from tripadvisor.com. The reviews were collected via web scraping, ensuring a diverse coverage across locations, and customer sentiments. We identified $8$ key aspects based on extensive consultations with experts in travel management and linguistics (Fig. 2). These aspects were chosen to reflect both operational and experiential dimensions of hotel service quality. Each review may correspond to multiple aspect labels, reflecting the inherently multi-label nature of user opinions. To ensure high-quality annotations, we employed $10$ trained human annotators for each review. For every review, the annotators assigned one or more aspect labels as applicable, and highlighted the exact sub-sentence or phrase responsible for each aspect label, thereby providing ground-truth supervision for interpretability evaluation. Prior to annotation, all annotators underwent training. To ensure annotation consistency and reliability, we implemented a quality control mechanism using in-context learning via GPT-4o (Dutta et al., 2025). Annotator performance was iteratively assessed by comparing human-labeled segments, enabling correction and refinement through active feedback loops. The resulting dataset consists of 3002 reviews, annotated at the sub-sentence level with multi-label aspect tags and text spans responsible for each label, making it one of the first of its kind, publicly available resources to support interpretability-grounded multi-aspect classification at such fine granularity. Brief statistical details of HR are presented in Table 1.
@@
Notably, sub-sentence level annotations are absent in the IMDb and TweetEVAL datasets.
Table 1. Statistics of proposed dataset: Hotel Reviews (HR)
@@
Attributes Values No. of Hotels & Locations 275 & 70 No. of Reviews & Sentences 3002 & 26519 No. of Aspects 8 Sentence per review (Min & Max) 1 & 100 Sentence per review (Avg. & Std. Dev.) 8.83 & 7.47 Word per review (Avg.) 102.39
07507501,5001{,}5002,2502{,}2503,0003{,} [^0] Guest ExperienceLocation & SurroundingHousekeepingRoom & AmenitiesFoodConciergePersonalizationPricing & Value2,4172{,}4171,9551{,}9551,8501{,}8502,1872{,}1872,2852{,}2852,2552{,}255976976900900Number of Reviews
Figure 2. Distribution of eight aspects in HR dataset
4.2. Evaluation Metrics
We evaluate model performance using task-specific metrics. For binary and multi-class classification, we report Accuracy ($\mathcal{A}$) and F1-score ($\mathcal{F}$). Aspect-wise evaluation additionally includes Precision ($\mathcal{P}r$) and Recall ($\mathcal{R}$). For multi-label classification, we use macro-averaged F1-score ($\mathcal{F}_{m}$) and Balanced Accuracy ($\mathcal{BA}_{m}$). We also compute Hamming Loss ($\mathcal{HL}$) and ${\mathcal{HL}}^{c}$ to quantify label-wise prediction accuracy; ${\mathcal{HL}}^{c}=(1-\mathcal{HL})\times 100\%$.
Table 2. Comparison with baseline and SOTA approaches

### v2

@@ III-B2 Dual Phase Learning
For sentence-level ground truth $\mathcal{G}_{2}$, we aggregate subsentence labels via $\mathcal{M}^{2}$. Document-level ground truth $\mathcal{G}_{3}\in\mathbb{R}^{1\times l}$ is directly taken from the dataset. It is worth noting that prototype and sentence-level labels are generated dynamically and depend on the current prototype configuration during the clustering phase.
III-B3 Prototype Projection and Label Classification
@@
Each prototype ${\mathcal{P}}_{i}\in\mathcal{P}$ is a high-dimensional vector residing in the same semantic space as subsentence embeddings. To ensure interpretability, we associate each prototype with its top-$K$ most similar subsentences from the training set, using cosine similarity. The top-$K$ nearest subsentences $\mathcal{T}_{K}({\mathcal{P}}_{i})$ are consolidated and passed to Gemini Pro 2.5 [gemini2023] to generate a concise summary reflecting their shared theme. This strategy ensures each prototype is grounded in semantically diverse and representative text snippets, offering robust explanations.
@@
Given a test document $\mathcal{D}$, we segment it into subsentences, and encode them using the LM-based encoder to get $\mathcal{E}\in\mathbb{R}^{t\times d}$. These embeddings are projected to the prototype space via the trained multi-head attention module to produce prototype-aware representations $\mathcal{H}$, which are then passed through the proto-to-label classifier $\tilde{\mathcal{G}}_{1}$. Sentence- and document-level predictions ($\tilde{\mathcal{G}}_{2}$, $\tilde{\mathcal{G}}_{3}$) are obtained via hierarchical mappings using $\mathcal{M}^{2}$ and $\mathcal{M}^{1}$. The final predicted class labels for the document are taken as $\tilde{\mathcal{G}}_{3}$. Simultaneously, we generate textual explanations by linking each subsentence to its closest prototype, and retrieving that prototype’s associated exemplars $\mathcal{T}_{K}({\mathcal{P}}_{i})$. This enables transparent reasoning behind predictions, making ProtoSiTex both accurate and interpretable.
@@
ProtoSiTex is best described as semi-interpretable because it combines prototype-based explanations with deep neural components. On the one hand, its adaptive prototypes provide human-understandable evidence by aligning subsentences with semantically coherent exemplars, allowing users to see why particular text spans support specific class predictions. On the other hand, the model relies on latent mechanisms, multi-head attention, embedding transformations, and hierarchical aggregation across subsentence, sentence, and document levels, which are not directly transparent to human reasoning. As a result, the explanations are faithful and intuitive at the prototype level but do not offer full interpretability of the entire decision process. This balance between interpretability and accuracy justifies labeling the model as semi-interpretable rather than fully interpretable.
III-C Training Strategy of ProtoSiTex
We adopt an alternate dual-phase training strategy, where each epoch comprises $i_{1}$ clustering and $i_{2}$ classification iterations. Prototype parameters are updated in the clustering phase (with classifiers frozen), and vice versa in the classification phase. This decoupling stabilizes training, prevents early overfitting, and allows prototypes to evolve independently before being used for label prediction. We now describe the loss functions associated with each phase.
III-C1 Loss Functions for Classification Phase
The classification phase is designed to facilitate hierarchical supervision by capturing semantic patterns from fine-grained (subsentence) to coarse-grained (document) levels. To train this phase, we use three loss components: (i) a prototype-to-label interpretability loss that reinforces the semantic alignment between prototypes and classes, (ii) a sentence-wise supervision loss that captures mid-level contextual cues, and (iii) a document-level loss leveraging the provided annotations. Each loss component is computed using binary cross entropy [PyTorch], and the combined loss is expressed as:
$\footnotesize\mathcal{L}_{cs}=\lambda_{1}\penalty 10000\ \mathcal{L}_{1}(\mathcal{G}_{1},\tilde{\mathcal{G}}_{1})+\lambda_{2}\penalty 10000\ \mathcal{L}_{2}(\mathcal{G}_{2},\tilde{\mathcal{G}}_{2})+\lambda_{3}\penalty 10000\ \mathcal{L}_{3}(\mathcal{G}_{3},\tilde{\mathcal{G}}_{3})$ (5)
where, $\mathcal{G}_{i}$ and $\tilde{\mathcal{G}}_{i}$ (for $i=1,2,3$) represent ground-truth and predicted labels at prototype, sentence, and document levels. The weights $\lambda_{i}$ control each loss term, with $\sum_{i=1}^{3}\lambda_{i}=1$. Subsentence and prototype-level annotations are optional; when unavailable, $\lambda_{1}$ and $\lambda_{2}$ can be reduced or set to zero, enabling the model to rely solely on document-level labels and support flexible supervision.
III-C2 Loss Functions for Clustering Phase
The clustering loss comprises two complementary components: proximity-based losses that enforce alignment between subsentence embeddings and prototypes, and structure regularization losses that ensure sparsity and diversity among the learned prototypes.
Proximity-based Losses
These losses promote semantic alignment and coverage. The prototype matching loss ensures that each subsentence $\mathcal{E}_{i}$ is closely aligned to its nearest prototype $\mathcal{P}_{j^{*}_{i}}$, encouraging relevance. Simultaneously, the interpretability loss ensures each prototype remains active by encouraging proximity to at least one subsentence, thus achieving coverage and preventing prototype collapse. The joint formulation is given by:
$\mathcal{L}_{prox}=-\frac{1}{b}\sum_{i=1}^{b}\left(\frac{\mathcal{E}_{i}^{\top}\mathcal{P}_{j^{*}_{i}}}{\lVert\mathcal{E}_{i}\rVert\lVert\mathcal{P}_{j^{*}_{i}}\rVert}\right)+\frac{1}{q}\sum_{j=1}^{q}\left(1-\max_{i=1}^{b}\left(\frac{\mathcal{E}_{i}^{\top}\mathcal{P}_{j}}{\lVert\mathcal{E}_{i}\rVert\lVert\mathcal{P}_{j}\rVert}\right)\right)$ (6)
where $j^{*}_{i}$ is the index of the prototype closest to subsentence $\mathcal{E}_{i}$ based on cosine similarity, and $b$ denotes the total number of subsentences across all reviews in a mini-batch.
Structure Regularization Losses
To enhance interpretability, we impose sparsity and diversity through regularization. The sparsity loss encourages selective activation, ensuring a limited set of subsentences influences each prototype. This reduces overlap and enhances clarity. The diversity loss, implemented via orthogonality regularization, pushes prototypes apart in embedding space, minimizing redundancy. The combined regularization is:
$\mathcal{L}_{reg}=-\frac{1}{q}\sum_{j=1}^{q}\log\left(\sum_{i=1}^{b}\frac{\exp(-\left\lVert\mathcal{E}_{i}-\mathcal{P}_{j}\right\rVert_{2})}{\sum_{k=1}^{q}\exp(-\left\lVert\mathcal{E}_{i}-\mathcal{P}_{k}\right\rVert_{2})}+\epsilon\right)$ (7)
$+\frac{1}{q(q-1)}\sum_{\begin{subarray}{c}i,j=1;\\ i\neq j\end{subarray}}^{q}\left(\frac{\mathcal{P}_{i}^{\top}\mathcal{P}_{j}}{\lVert\mathcal{P}_{i}\rVert\lVert\mathcal{P}_{j}\rVert}\right)^{2}$
where, the first term encourages peaked assignment distributions (sparsity), and the second term enforces pairwise dissimilarity (diversity) between prototypes.
To enforce semantic alignment with classes, we augment the clustering loss with the supervised classification loss $\mathcal{L}_{cs}$. The final clustering-phase loss ($\mathcal{L}_{cr}$) is:
$\footnotesize\mathcal{L}_{cr}=\alpha_{1}\penalty 10000\ \mathcal{L}_{prox}+\alpha_{2}\penalty 10000\ \mathcal{L}_{reg}+\alpha_{3}\penalty 10000\ \mathcal{L}_{cs}$ (8)
where, $\alpha_{1},\alpha_{2},\alpha_{3}$ are tunable weights summing to 1, controlling the influence of each component.
The classification phase refines prototypes into interpretable classes, enhancing clustering and enabling generalizable, multi-granular predictions.
The computational complexity of ProtoSiTex is analyzed in Appendix A of the supplementary file. Appendix B provides a detailed discussion on the stability of prototype updates and the influence of loss-weight parameters on model behavior.
IV Experiments
This section describes the datasets employed and reports the corresponding experimental results.
IV-A Datasets
@@
To evaluate the effectiveness of ProtoSiTex and compare it against state-of-the-art (SOTA) methods, we conducted comprehensive experiments on benchmark datasets IMDb [IMDb], TweetEVAL [TweetEVAL], and our newly proposed hotel reviews dataset (HR). These datasets encompass a diverse range of classification tasks: binary, multi-class, and multi-label, respectively.
@@
(i) IMDb [IMDb]: Consisting of 50K movie reviews for binary sentiment classification (positive/ negative).
@@
(ii) TweetEVAL [TweetEVAL]: A multiclass dataset comprising 5052 tweets, with each tweet annotated with one of four emotion labels: anger, joy, optimism, or sadness.
07507501,5001{,}5002,2502{,}2503,0003{,} [^0] Room & AmenitiesPricing & ValuePersonalizationLocation & SurroundingHousekeepingGuest ExperienceFoodConcierge2,1872{,}1879009009769761,9551{,}9551,8501{,}8502,4172{,}4172,2852{,}2852,2552{,}255Document-level review count
Figure 2: Distribution of eight aspects in HR
@@
(iii) Proposed Dataset on Hotel Reviews (HR): To support fine-grained, interpretable, multi-aspect classification in the hospitality domain, we introduce a novel dataset comprising hotel reviews, curated from tripadvisor.com. The reviews were collected via web scraping, ensuring a diverse coverage across locations, and customer sentiments. We identified $8$ key aspects based on extensive consultations with experts in travel management and linguistics (Fig. 2). These aspects were chosen to reflect both operational and experiential dimensions of hotel service quality. Each review may correspond to multiple aspect labels, reflecting the inherently multi-label nature of user opinions. To ensure high-quality annotations, we employed $10$ trained human annotators for each review. For every review, the annotators assigned one or more aspect labels as applicable, and highlighted the exact subsentence or phrase responsible for each aspect label, thereby providing ground-truth supervision for interpretability evaluation. Prior to annotation, all annotators underwent training. We implemented a quality control mechanism using in-context learning via GPT-4o [ARTICLE_AAAI_2025]. Annotator performance was iteratively assessed by comparing human-labeled segments, enabling correction and refinement through active feedback loops. To ensure annotation consistency and reliability, we quantified inter-annotator agreement using Krippendorff’s alpha ($\upalpha$) [krippendorff1970]. The obtained $\upalpha=0.85$ indicates strong inter-annotator agreement, validating the annotation procedure. The resulting dataset consists of 3002 document-level and 26519 sentence-level reviews, annotated at the subsentence level with multi-label aspect tags and text spans responsible for each label, making it one of the first of its kind to support interpretability-grounded multi-aspect classification at such fine granularity. Brief statistical details of HR are presented in Table I.
@@
Notably, subsentence level annotations are absent in the IMDb and TweetEVAL datasets.
IV-B Experimental Setup
All experiments were conducted on a Dell Precision 5860 Tower equipped with PyTorch 2.5.1, CUDA 12.2, an NVIDIA RTX A6000 GPU with 48 GB VRAM, and 256 GB RAM. The HR dataset was split into training, validation, and test sets in 7:2:17{:}2{:}1. For IMDb [IMDb] and TweetEVAL [TweetEVAL], we used the given training/ validation/ test sets. Each experiment was run five times, and the mean test performance is reported here.
Table II presents the hyperparameters engaged in ProtoSiTex. To obtain high-dimensional, contextually rich text embeddings, we incorporate a language model, called RoBERTa [liu2019roberta]. Specifically, it is a robustly optimized BERT variant, which we employed in its large-weight version pretrained on large English-language internet corpora using the masked language modeling objective; RoBERTa-Large remains tunable within our framework. We maintain a consistent dropout rate of 0.1 across all trainable weights. To avoid overfitting, early stopping with a patience of three epochs based on validation performance is applied. Additionally, we set a maximum training limit of 100 epochs. Each epoch includes 10 ($i_{1}$) iterations of the clustering phase and 100 ($i_{2}$) iterations of the classification phase, ensuring a balanced optimization of both phases.
TABLE I: Statistics of proposed dataset: Hotel Reviews (HR)
@@ IV-B Experimental Setup
Attributes Values No. of Hotels & Locations 275 & 70 No. of Reviews & Sentences 3002 & 27074 No. of Aspects 8 Sentences per review (min & max) 1 & 92 Sentences per review (mean $\pm$ stdev) 9.01 $\pm$ 7.69 Words per review (mean) 120.82
TABLE II: Brief details of hyperparameters
Attribute Value Optimizer AdamW Learning rate (Clustering-phase) $2\times 10^{-5}$ Learning rate (Classification-phase) $2\times 10^{-5}$ Batch size 64 LM encoder: Feature dimension 1024 No. of dense layer 2 Dimension of dense layer 256 & 64 Clustering loss parameters ($\alpha_{1}$, $\alpha_{2}$, $\alpha_{3}$ ) [0-1] Classification loss parameters ($\lambda_{1}$, $\lambda_{2}$, $\lambda_{3}$) [0-1] No. of attention heads ($h$) [1-32] No. of learnable prototypes ($q$) [4-64]
TABLE III: Comparison with baseline and SOTA approaches
@@
IV-C Evaluation Metrics
We evaluate model performance using task-specific metrics. For binary and multi-class classification, we report Accuracy ($\mathcal{A}$) and F1-score ($\mathcal{F}$); aspect-wise evaluation additionally includes Precision ($\mathcal{P}r$) and Recall ($\mathcal{R}$); for multi-label classification, we use macro-averaged F1-score ($\mathcal{F}_{m}$) and Balanced Accuracy ($\mathcal{BA}_{m}$); all are in percentage. We also compute Hamming loss ($\mathcal{HL}$) to analyze label-wise prediction performance, and report complementary Hamming loss ${\mathcal{HL}}^{c}$ here; ${\mathcal{HL}}^{c}=(1-\mathcal{HL})\times 100\%$.
IV-D Results & Comparative Study
To compare with ProtoSiTex, we chose 6 SOTA prototypical learning models: ProSeNet [10.1145/3292500.3330908-ProsNet], ProtoTEX [das-etal-2022-prototex], ProtoryNet [10.5555/3648699.3648963-ProtoryNet], Proto-lm [xie-etal-2023-proto], GAProtoNet [wen-etal-2025-gaprotonet], and ProtoLens [protolens]. In addition to this, we selected 9 black-box LMs: ALBERT [ALBERT], DeBERTaV3 [DeBERTaV3] BERT [devlin-etal-2019-bert], DistilBERT [sanh2019distilbert], ELECTRA [ELECTRA], XLNET [yang2019xlnet], ModernBERT [ModernBERT], RoBERTa and RoBERTa-Large [liu2019roberta], and 2 text classification methods: TextGCN [textgcn] and BertGCN [bertgcn], without incorporating interpretability.
In Table III, we present the comparative study. This study reveals that existing prototype-based SOTA approaches often sacrifice performance for interpretability, leading to significantly lower results compared to baselines, particularly on challenging multi-label datasets such as HR. In contrast, ProtoSiTex successfully overcame this limitation by narrowing the performance gap with black-box LMs, while preserving inherent interpretability across varying dataset complexities. Specifically, ProtoSiTex achieved improvements of 0.04% and 2.85% in $\mathcal{A}$} on the IMDb and TweetEVAL datasets, respectively, and delivered a substantial 9.98 % improvement in $\mathcal{F}_{m}$} on HR, compared to the best prior prototype-based methods. These results across three diverse tasks highlight ProtoSiTex’s generalizability and applicability.
${\mathcal{F}}_{m}$${\mathcal{BA}}_{m}$${\mathcal{HL}}^{c}$75758080858590909595K-meansKaimingXavierRandom
${\mathcal{F}}_{m}$${\mathcal{BA}}_{m}$${\mathcal{HL}}^{c}$6060707080809090100100w/o ATProtoSiTex
${\mathcal{F}}_{m}$${\mathcal{BA}}_{m}$${\mathcal{HL}}^{c}$2020404060608080100100Euclidean Dist.Cosine Sim.ProtoSiTex
[…]
Figure 3: Performance analysis of ProtoSiTex on HR: (a) with various prototype initializers; (b) without alternate training (w/o AT); (c) across different similarity measures; with a varying number of (d) prototypes, and (e) heads
[…]
Figure 4: Impact of hyperparameters $\alpha_{i}$ and $\lambda_{i}$ on HR
IV-E Impact of Model Components
We here report the impact of modules of ProtoSiTex on HR dataset only due to space constraints, which is representative enough to illustrate the observed trends.
IV-E1 Impact of Different Prototype Initializers
We evaluated four prototype initialization strategies: random, K-means centroids, Kaiming, and Xavier. As shown in Fig. 3(a), random initialization consistently outperformed the others across $\mathcal{F}_{m}$, $\mathcal{BA}_{m}$, and ${\mathcal{HL}}^{c}$. This performance gain may be attributed to the increased diversity introduced early in the training, which helps avoid premature convergence and overfitting.
IV-E2 Impact of Alternate Training
ProtoSiTex adopts a dual-phase training strategy, alternating between unsupervised prototype learning and supervised classification. To evaluate its efficacy, we compared model variants trained with and without alternation. As shown in Fig. 3(b), alternate training significantly improved performance by enabling the model to discover semantically meaningful prototype clusters, and align them with downstream classification objectives. In contrast, without alternation, the model struggled to learn task-specific prototypes, lowering performance.
IV-E3 Impact of Similarity Function
IV-F Impact of Hyperparameters
ProtoSiTex includes several key hyperparameters: the number of prototypes ($q$), number of heads ($h$), loss weights for clustering ($\alpha_{1}$, $\alpha_{2}$, $\alpha_{3}$), and classification ($\lambda_{1}$, $\lambda_{2}$, $\lambda_{3}$). To systematically investigate the influence of these hyperparameters on model performance, we conducted a series of experiments on the HR. As shown in Fig. 3(d), increasing $q$ initially boosts performance by capturing diverse semantic concepts, but excessive values lead to redundancy and over-smoothing. A similar trend is observed for $h$ in Fig. 3(e). Furthermore, Fig. 4 shows that assigning $\alpha_{1}=0.01$, $\alpha_{2}=0.98$, $\alpha_{3}=0.01$, and $\lambda_{1}=0.6$, $\lambda_{2}=0.2$, $\lambda_{3}=0.2$ yields optimal results. This multi-level supervision prevented the model from overfitting to a single level of granularity, supporting both local and global interpretability and accuracy.
[…]
Figure 5: Aspect-wise performance of ProtoSiTex on HR